# Generalizable LLM Learning of Graph Synthetic Data with Reinforcement Learning

### v2

@@ 5 Analysis
While using purely synthetic data and designed rewards to tune the model using RL achieves good results for certain tasks, the RL methods we use (GRPO and DPO) still cannot achieve a universal performance boost when testing on different domains. We further conduct additional analysis using Qwen2.5-7B-Instruct to understand the gap between synthetic and real-world reasoning.
@@ 5.1 Mixing Real-World Data
While the original goal for this work is to understand and evaluate if RL can further generalize synthetic graph reasoning to real-world implicit graph reasoning tasks, we believe that normally we can acquire a small amount of labelled data for a real-world SFT stage with human annotation. We simulate this phase by introducing an additional SFT training using certain amount of labelled real-world data, and then we perform the RL stage where all data instances and rewards are synthetic. While the SFT stage somehow provides the model clues about how to answer each kind of question, we are interested in finding out if further synthetic RL training can increase the model’s performance.
We derive a total of 300/1200/6000 samples with question and ground-truth pairs from the following six datasets: StrategyQA, Knowledge Crosswords, WikiDYK (Reliability), WikiDYK (Factual), ExplaGraphs, and Proscript, to simulate a small-sample SFT stage. Each real-world SFT set consists of $1/6$ of the total data. We train models using SFT for 3 epochs, and then continue to use synthetic graph tasks to GRPO tune the model using process-based reward. Results are shown in Table 3.
Qwen2.5-7B-Instruct StrategyQA K-C WikiDYK-R WikiDYK-F ExplaGraphs Proscript Zero-Shot 0.702 0.504 0.067 0.431 0.829 0.592 Synthetic SFT 0.615 0.268 0.069 0.531 0.668 0.488 Synthetic SFT/RL Best 0.717 0.523 0.070 0.535 0.870 0.605 Real-world SFT 6k 0.682 0.785 0.096 0.676 0.932 0.851 /w synthetic RL 0.671 0.791 0.092 0.676 0.936 0.789 Real-world SFT 1.2k 0.698 0.654 0.077 0.660 0.915 0.795 /w synthetic RL 0.661 0.645 0.072 0.648 0.922 0.757 Real-world SFT 0.3k 0.708 0.579 0.069 0.601 0.901 0.729 /w synthetic RL 0.742 0.449 0.066 0.641 0.907 0.638
Out of 18 RL results, 6 achieved improvements w.r.t. their real-world SFT setting and other baselines; however, none of the improvements is statistically significant. Also, RL’s effects are still task-dependent, without bringing a universal improvement across all real-world tasks. In general, under the current setting, even mixing a small to medium portion of real-world data, a synthetic RL stage after that is not the golden solution to generalize reasoning beyond synthetic patterns.
@@ 5.2 Translation Between Single-Step and Multi-Step
As shown in Table 2, results on StrategyQA and Knowledge Crosswords after RL training (with or without the optional SFT stage) do not show any significant performance increase. Intuitively, single-step results should generally be better, or at least not worse, than multi-step results due to the limitation of LLM’s multi-step reasoning capabilities. We are interested in whether previous results indicate a limitation in knowledge, a reasoning gap between single-step and multi-step implicit graph reasoning, or any other interesting findings. We build upon two datasets and extract a group of single-step questions from each multi-step question to probe the synthetic-trained language model’s response. Results are shown in Figure 2. Examples of single-step questions for two datasets and implementation details, and additional analysis are presented in Appendix E.
@@
We see out of all the correct single-step questions, there are at least 25%, or even as much as 46%, of the incorrect multi-step results across all training settings. This means that there is indeed a gap between single-step and multi-step reasoning, and even using synthetic RL cannot solve this problem. Another interesting finding is that, out of all the correct multi-step results, at least half of the correct results come from a not-completely-correct single-step results. For instance, the model successfully answers a multi-step question ‘A->C’ which needs the information of single-step ‘A->B’ and ‘B->C’, but fails to answer at least one of the single-step questions. Also, the ratio of incorrect single-step out of all correct multi-step answers tends to increase with more real-world SFT data. This suggests that either the model is hallucinating the final answer, or the model is using alternative methods to solve the multi-step question. Both scenarios (using wrong/unexplainable intermediate steps to solve a multi-step question, or failing to get the correct final result even with correct intermediate steps) are not desired in this multi-step reasoning setting.
@@
In summary, while knowledge gap may be a partial reason for the incapability of generalizing to real-world tasks, two more significant caveats arise. First, RL struggles to patch the compositionality gap from single steps to the full problem. Second, performance increases, if any, are more often caused by multi-step hallucination, as models might provide correct answers to the full problem without an accurate understanding of the underlying intermediate steps. These two limitations highlight the caveats of RL learning with synthetic graph data and motivate solutions as future work.
5.3 Correlation Analysis
We investigate the correlation between various performance and reward metrics. First, to better understand the reward design’s effect, we analyze the correlation between the synthetic task’s accuracy performance and the reward score achieved by the model. The Pearson correlation coefficient is 0.951: our reward design is highly correlated with the performance of the model, proving the reward function’s effectiveness. Second, we analyze the correlation between synthetic performance and real-world performance (denoted by the mean performance of all real-world tasks). The correlation is -0.336, which aligns with our finding that performance gains provided by synthetic training are mixed across models and different tasks. Detailed results are shown in Figure 3.
@@ 5.3 Correlation Analysis
6 Related Work
Reinforcement Learning with LLMs
Large language models (LLMs) have been increasingly fine-tuned with reinforcement learning from human feedback (RLHF) to improve alignment, safety, and reasoning [55, 5, 36]. Early applications of RLHF demonstrated significant gains in complex tasks such as text summarization and instruction-following [72, 55] compared to previous models[6]. To reduce the cost of human feedback, RLAIF uses AI-generated preferences to train reward models with comparable performance [41]. Most RL on LLM pipelines optimize LLM policies with on-policy algorithms, notably Proximal Policy Optimization (PPO) [65], which originates from TRPO [64], to iteratively maximize reward model outputs while constraining divergence from the base model. Group Relative Policy Optimization (GRPO) eliminates the need for a separate value network by normalizing rewards across batches, greatly improving training efficiency on reasoning tasks [67]. On the other hand, Direct Preference Optimization (DPO), as an off-policy algorithm, reframes preference alignment as a simple supervised objective without costly sampling [60]. RL methods have substantially advanced the multi-step reasoning capabilities of LLMs [12, 23, 43, 40, 33, 49, 37], with researchers exploring more capabilities of RL [98, 104, 68, 82, 86, 9].
Graph Reasoning with LLMs
Recent work has extended LLMs to reason over graph-structured data by encoding explicit graph structures, fine-tuning with graph instructions, and retrieval augmentation. A key idea is to represent graphs in an LM-readable form or code, either by linearizing graph topology into text prompts or by injecting edge lists and paths into the context [19, 84, 28, 51]. Beyond prompting, specialized graph instruction tuning has emerged: LLMs are fine-tuned on graph reasoning tasks with potential diverse modality to better internalize structured knowledge [10, 50, 57, 85, 44, 14, 75, 30, 102, 83, 16, 7], or further enhanced with graph structures [46, 8, 88]. Such models outperform zero-shot LLMs on tasks like knowledge graph question answering and multi-hop reasoning, demonstrating that integrating graph context can curb hallucinations and improve relational inference [29, 25]. However, recent benchmarks suggest that while fine-tuning on synthetic graph data can teach LLMs specific patterns, these models often cannot fully transfer [102, 12, 76] or struggle to transfer beyond their training distribution or generalize to real-world tasks [100, 25]. This gap has spurred new efforts to improve LLMs’ graph reasoning robustness, though achieving reliable out-of-distribution generalization remains an open challenge.
7 Conclusion
In this work, we investigate using RL to generalize LLM graph learning beyond synthetic problems. We use synthetic graph tasks including connectivity and shortest path, and implement the RL reward using rule-based rewards with two designs: process-based and solution-based reward, finding that process-based reward consistently outperforms solution-based reward. While models purely trained on synthetic problems with RL can lead to overall synthetic performance increases and partial performance increases on real-world tasks, RL on synthetic data does not provide a universal solution to all real-world tasks. Further, we analyze the performance bottleneck of current LLMs graph reasoning, including two important caveats: failure to generalize from single-step to multi-step reasoning, and potential hallucination from single-step reasoning to multi-step reasoning. With the partial success of synthetic RL on real-world tasks, we believe further research is in need to fully understand the compositionality, reliability, and explainability of graph-related single-step to multi-step generalization of LLMs.
Appendix A Limitations
Adopted Methods
We only adopt two representative RL methods (GRPO and DPO) on LLMs, while there are still a wide range of on-policy optimization methods [99, 65, 97, 95, 26, 61, 4] and off-policy optimization methods [92, 18, 53] with possible model-based rewards [79]. With that in mind, in general, we do see that different methods are enhancing training efficiency and stability and should not impact the overall results of our work greatly.
We also notice there are different methods besides RL, aiming to improve general reasoning capabilities of LLMs, including test-time scaling [54, 70, 66, 43, 104, 34, 96] and sub-response level reward or supervision [91, 79, 45]. We leave experimenting using these methods for future work.
Adopted Models
We only adopt two representative open-source models (Qwen2.5 and LLaMA3.1 with relative small size (7B and 8B respectively) [42]. While there are still a wide range of open-source LLMs [22, 93, 38, 47, 77] and close-source LLMs [1, 24, 2, 3, 36, 35, 58] with different model sizes, we do believe the ultimate goal is to achieve a universal and general good performance across all models and all sizes, especially with potential of fully utilizing synthetic data on real-world tasks, and reliable and huallicination-free reasoning steps in multi-step reasoning tasks.
Real-World Datasets
We only adopt 8 real-world tasks within 3 main categories (multi-hop QA, commonsense reasoning, and action planning). There are still plenty of datasets which belongs to these three categories [52, 15, 74, 21, 94, 48, 11] and other reasoning related datasets and domains [31, 13, 78, 59, 39, 103, 73, 90, 71, 89, 87]. We also notice that our selected datasets may not thoroughly represent real-world datasets and tasks perfectly, but in general they still have a large gap with purely synthetic data generated by algorithms and symbolic representations.
@@ Appendix B Reward Implementation Details
  • Correct answer reward: $r_{\text{correct\_answer}}=+1$
  • Incorrect answer penalty: $r_{\text{incorrect\_answer\_penalty}}=0$
  • Hallucination penalty: $r_{\text{hallucination\_penalty}}=-2$
  • Correct reasoning step reward: $r_{\text{correct\_step}}=+0.05$
  • Incorrect reasoning penalty: $r_{\text{incorrect\_step\_penalty}}=0$
  • Correct format reward: $r_{\text{format\_reward}}=+0.1$
  • Format error penalty: $r_{\text{format\_penalty}}=0$
@@
$A$ and $B$: two nodes in the given graph.
@@
ground_truth: ground truth connectivity (“yes” if $A$ and $B$ is connected, “no” otherwise).
@@
$r\leftarrow 0$triangleright initialize reward score
@@
     $r\leftarrow r+r_{\text{format\_penalty}}$triangleright penalize formatting issue
@@
     $r\leftarrow r+r_{\text{format\_reward}}$triangleright reward correct format
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
     $r\leftarrow r+r_{\text{correct\_answer}}$triangleright reward correct conclusion
@@
     $r\leftarrow r+r_{\text{incorrect\_answer\_penalty}}$triangleright penalize incorrect conclusion
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v3

@@ 5 Analysis
While using purely synthetic data and designed rewards to tune the model using post-training alignment achieves good results for certain tasks, the alignment methods we use (GRPO and DPO) still cannot achieve a universal performance boost when testing on different domains. We further conduct additional analysis using Qwen2.5-7B-Instruct to understand the gap between synthetic and real-world reasoning.
@@ 5.1 Mixing Real-World Data
While the original goal for this work is to understand and evaluate if alignment can further generalize synthetic graph reasoning to real-world implicit graph reasoning tasks, we believe that normally we can acquire a small amount of labelled data for a real-world SFT stage with human annotation. We simulate this phase by introducing an additional SFT training using certain amount of labelled real-world data, and then we perform the alignment stage where all data instances and rewards are synthetic. While the SFT stage somehow provides the model clues about how to answer each kind of question, we are interested in finding out if further synthetic alignment training can increase the model’s performance.
We derive a total of 300/1200/6000 samples with question and ground-truth pairs from the following six datasets: StrategyQA, Knowledge Crosswords, WikiDYK (Reliability), WikiDYK (Factual), ExplaGraphs, and Proscript, to simulate a small-sample SFT stage. Each real-world SFT set consists of $1/6$ of the total data. We train models using SFT for 3 epochs, and then continue to use synthetic graph tasks to GRPO tune the model using process-based reward. Results are shown in Appendix B Table 3.
Out of 18 alignment results, 6 achieved improvements w.r.t. their real-world SFT setting and other baselines; however, none of the improvements is statistically significant. Also, alignment’s effects are still task-dependent, without bringing a universal improvement across all real-world tasks. In general, under the current setting, even mixing a small to medium portion of real-world data, a synthetic alignment stage after that is not the golden solution to generalize reasoning beyond synthetic patterns.
@@ 5.2 Translation Between Single-Step and Multi-Step
As shown in Table 2, results on StrategyQA and Knowledge Crosswords after alignment training (with or without the optional SFT stage) do not show any significant performance increase. Intuitively, single-step results should generally be better, or at least not worse, than multi-step results due to the limitation of LLM’s multi-step reasoning capabilities. We are interested in whether previous results indicate a limitation in knowledge, a reasoning gap between single-step and multi-step implicit graph reasoning, or any other interesting findings. We build upon two datasets and extract a group of single-step questions from each multi-step question to probe the synthetic-trained language model’s response. Results are shown in Appendix B Figure 2. Examples of single-step questions for two datasets and implementation details, and additional analysis are presented in Appendix E.
@@
We see out of all the correct single-step questions, there are at least 25%, or even as much as 46%, of the incorrect multi-step results across all training settings. This means that there is indeed a gap between single-step and multi-step reasoning, and even using synthetic alignment cannot solve this problem. Another interesting finding is that, out of all the correct multi-step results, at least half of the correct results come from a not-completely-correct single-step results. For instance, the model successfully answers a multi-step question ‘A->C’ which needs the information of single-step ‘A->B’ and ‘B->C’, but fails to answer at least one of the single-step questions. Also, the ratio of incorrect single-step out of all correct multi-step answers tends to increase with more real-world SFT data. This suggests that either the model is hallucinating the final answer, or the model is using alternative methods to solve the multi-step question. Both scenarios (using wrong/unexplainable intermediate steps to solve a multi-step question, or failing to get the correct final result even with correct intermediate steps) are not desired in this multi-step reasoning setting.
@@
In summary, while knowledge gap may be a partial reason for the incapability of generalizing to real-world tasks, two more significant caveats arise. First, alignment struggles to patch the compositionality gap from single steps to the full problem. Second, performance increases, if any, are more often caused by multi-step hallucination, as models might provide correct answers to the full problem without an accurate understanding of the underlying intermediate steps. These two limitations highlight the caveats of alignment learning with synthetic graph data and motivate solutions as future work.
6 Related Work
Post-training with LLMs
Large language models (LLMs) have been increasingly fine-tuned with reinforcement learning from human feedback (RLHF) to improve alignment, safety, and reasoning Ouyang et al. (2022); Bai et al. (2022); Jaech et al. (2024). Early applications of RLHF demonstrated significant gains in complex tasks such as text summarization and instruction-following (Stiennon et al., 2022; Ouyang et al., 2022) compared to previous models(Brown et al., 2020). To reduce the cost of human feedback, RLAIF uses AI-generated preferences to train reward models with comparable performance (Lee et al., 2023). Most RL on LLM pipelines optimize LLM policies with on-policy algorithms, notably Proximal Policy Optimization (PPO) (Schulman et al., 2017), which originates from TRPO (Schulman et al., 2015), to iteratively maximize reward model outputs while constraining divergence from the base model. Group Relative Policy Optimization (GRPO) eliminates the need for a separate value network by normalizing rewards across batches, greatly improving training efficiency on reasoning tasks (Shao et al., 2024). On the other hand, Direct Preference Optimization (DPO), reframes preference alignment as a simple supervised objective without costly sampling (Rafailov et al., 2023). Alignment methods have substantially advanced the multi-step reasoning capabilities of LLMs (Chu et al., 2025; Guo et al., 2025a; Li et al., 2025; Kumar et al., 2024; Hu, 2025; Luo et al., 2025; Jain et al., 2025), with researchers exploring more capabilities of post-training alignment (Yue et al., 2025; Zuo et al., 2025; Shen et al., 2025; Wan et al., 2025; Wei et al., 2025; Chen et al., 2025).
Graph Reasoning with LLMs
Recent work has extended LLMs to reason over graph-structured data by encoding explicit graph structures, fine-tuning with graph instructions, and retrieval augmentation. A key idea is to represent graphs in an LM-readable form or code, either by linearizing graph topology into text prompts or by injecting edge lists and paths into the context (Fatemi et al., 2024; Wang et al., 2023; Han et al., 2024; Madaan et al., 2022). Beyond prompting, specialized graph instruction tuning has emerged: LLMs are fine-tuned on graph reasoning tasks with potential diverse modality to better internalize structured knowledge (Chen et al., 2024c; LUO et al., 2024; Perozzi et al., 2024; Wang et al., 2024b; Li et al., 2024b; Das et al., 2024; Tang et al., 2024; He et al., 2024b; Zhu et al., 2024; Wang et al., 2024a; Deng et al., 2024; Chen et al., 2024a), or further enhanced with graph structures (Lin et al., 2024; Chen et al., 2024b; Wu et al., 2024). Such models outperform zero-shot LLMs on tasks like knowledge graph question answering and multi-hop reasoning, demonstrating that integrating graph context can curb hallucinations and improve relational inference (He et al., 2024a; Guo et al., 2023). However, recent benchmarks suggest that while fine-tuning on synthetic graph data can teach LLMs specific patterns, these models often cannot fully transfer (Zhu et al., 2024; Chu et al., 2025; Tang et al., 2025) or struggle to transfer beyond their training distribution or generalize to real-world tasks (Zhang et al., 2024b; Guo et al., 2023). This gap has spurred new efforts to improve LLMs’ graph reasoning robustness, though achieving reliable out-of-distribution generalization remains an open challenge.
7 Conclusion
In this work, we investigate using post-training alignment to generalize LLM graph learning beyond synthetic problems. We use synthetic graph tasks including connectivity and shortest path, and implement the alignment reward using rule-based rewards with two designs: process-based and solution-based reward, finding that process-based reward consistently outperforms solution-based reward. While models purely trained on synthetic problems with alignment can lead to overall synthetic performance increases and partial performance increases on real-world tasks, post-training alignment on synthetic data does not provide a universal solution to all real-world tasks. Further, we analyze the performance bottleneck of current LLMs graph reasoning, including two important caveats: failure to generalize from single-step to multi-step reasoning, and potential hallucination from single-step reasoning to multi-step reasoning. With the partial success and important findings of synthetic alignment on real-world tasks, we believe further research is in need to fully understand the compositionality and explainability of graph-related generalization of LLMs.
Limitations
Adopted Methods
We only adopt two representative post-training alignment methods (GRPO and DPO) on LLMs, while there are still a wide range of on-policy optimization methods (Zhang et al., 2024a; Schulman et al., 2017; Yuan et al., 2025b; Yu et al., 2025; Guo et al., 2024; Rosset et al., 2024; Agarwal et al., 2024) and off-policy optimization methods (Xu et al., 2024; Ethayarajh et al., 2024; Meng et al., 2024) with possible model-based rewards (Uesato et al., 2022). With that in mind, in general, we do see that different methods are enhancing training efficiency and stability and should not impact the overall results of our work greatly.
We also notice there are different methods besides adopted post-training alignment, aiming to improve general reasoning capabilities of LLMs, including test-time scaling (Muennighoff et al., 2025; Snell et al., 2024; Setlur et al., 2025; Li et al., 2025; Zuo et al., 2025; Huang et al., 2025; Yuan et al., 2025a) and sub-response level reward or supervision (Xiong et al., 2024; Uesato et al., 2022; Lightman et al., 2023). We leave experimenting using these methods for future work.
Adopted Models
We only adopt two representative open-source models (Qwen2.5 and LLaMA3.1 with relative small size (7B and 8B respectively) (Li et al., 2024a). While there are still a wide range of open-source LLMs (Grattafiori et al., 2024; Yang et al., 2024; Jiang et al., 2023; Liu et al., 2024a; Team et al., 2024) and close-source LLMs (ant, ; Guo et al., 2025b; dee, ; ope, ; Jaech et al., 2024; Hurst et al., 2024; Pichai, 2024) with different model sizes, we do believe the ultimate goal is to achieve a universal and general good performance across all models and all sizes, especially with potential of fully utilizing synthetic data on real-world tasks, and reliable and huallicination-free reasoning steps in multi-step reasoning tasks.
Real-World Datasets
We only adopt 8 real-world tasks within 3 main categories (multi-hop QA, commonsense reasoning, and action planning). There are still plenty of datasets that belong to these three categories (Mavi et al., 2024; Davis, 2023; Talmor et al., 2021; Ghazal et al., 2013; Yang et al., 2025; Liu et al., 2024b; Choi et al., 2024) and other reasoning related datasets and domains (Hendrycks et al., 2021; Cobbe et al., 2021; Tong et al., 2024; Qiu et al., 2025; Jimenez et al., 2023; Zhuo et al., 2024; Sui et al., 2024; Xiong et al., 2023; Sprague et al., 2023; Xin et al., 2024; Wen et al., 2024). We also notice that our selected datasets may not thoroughly represent real-world datasets and tasks perfectly, but in general they still have a large gap with purely synthetic data generated by algorithms and symbolic representations.
Appendix A Tables and Graphs
Table for Section 5.1 of results of mixing real-world data, figure for Section 5.2 of comparing between multi-hop and single-hop and figure for of correlation analysis are attached here.
Qwen2.5-7B-Instruct StrategyQA K-C WikiDYK-R WikiDYK-F ExplaGraphs Proscript Zero-Shot 0.702 0.504 0.067 0.431 0.829 0.592 Synthetic SFT 0.615 0.268 0.069 0.531 0.668 0.488 Synthetic SFT/Alignment Best 0.717 0.523 0.070 0.535 0.870 0.605 Real-world SFT 6k 0.682 0.785 0.096 0.676 0.932 0.851 /w synthetic Align 0.671 0.791 0.092 0.676 0.936 0.789 Real-world SFT 1.2k 0.698 0.654 0.077 0.660 0.915 0.795 /w synthetic Align 0.661 0.645 0.072 0.648 0.922 0.757 Real-world SFT 0.3k 0.708 0.579 0.069 0.601 0.901 0.729 /w synthetic Align 0.742 0.449 0.066 0.641 0.907 0.638
@@ Appendix B Reward Implementation Details
  • Correct answer reward:
    $r_{\text{correct\_answer}}=+1$
  • Incorrect answer penalty:
    $r_{\text{incorrect\_answer\_penalty}}=0$
  • Hallucination penalty:
    $r_{\text{hallucination\_penalty}}=-2$
  • Correct reasoning step reward:
    $r_{\text{correct\_step}}=+0.05$
  • Incorrect reasoning penalty:
    $r_{\text{incorrect\_step\_penalty}}=0$
  • Correct format reward:
    $r_{\text{format\_reward}}=+0.1$
  • Format error penalty:
    $r_{\text{format\_penalty}}=0$
@@
AA and $B$: two nodes in the given graph.
@@
ground_truth: ground truth connectivity (“yes” if AA and $B$ is connected, “no” otherwise).
@@
$r\leftarrow 0$ $\triangleright$ initialize reward score
@@
  $r\leftarrow r+r_{\text{format\_penalty}}$ $\triangleright$ penalize formatting issue
@@
  $r\leftarrow r+r_{\text{format\_reward}}$ $\triangleright$ reward correct format
@@
$(\textit{thought},\textit{reasoningProcess},\textit{answer})\leftarrow\textsc{SeparateResponse}(R)$
@@
if $\textit{type}=\text{``process''}$  then
  $\mathit{rewards}\leftarrow[\,]$ $\triangleright$ initialize an empty list
@@
     Append $r_{\text{hallucination\_penalty}}$ to $\mathit{rewards}$ $\triangleright$ penalize hallucinated graph elements
@@
     Append $r_{\text{correct\_step}}$ to $\mathit{rewards}$ $\triangleright$ reward a correct reasoning step
@@
     Append $r_{\text{incorrect\_step\_penalty}}$ to $\mathit{rewards}$ $\triangleright$ penalize a step with incorrect format
@@
  $r_{\text{avg}}\leftarrow\text{Average}(\mathit{rewards})$ $\triangleright$ average the process reward
@@
  $r\leftarrow r+r_{\text{correct\_answer}}$ $\triangleright$ reward correct conclusion
@@
  $r\leftarrow r+r_{\text{incorrect\_answer\_penalty}}$ $\triangleright$ penalize incorrect conclusion
@@
$(\textit{thought},\textit{reasoningProcess},\textit{answer})\leftarrow\textsc{SeparateResponse}(R)$
if $\textit{type}=\text{``process''}$ then
  $\mathit{rewards}\leftarrow[\,]$
@@
     Append $r_{\text{hallucination\_penalty}}$ to $\mathit{rewards}$
@@
     Append $r_{\text{correct\_step}}$ to $\mathit{rewards}$
@@
     Append $r_{\text{incorrect\_step\_penalty}}$ to $\mathit{rewards}$
@@
  $r_{\text{avg}}\leftarrow\text{Average}(\mathit{rewards})$
@@
$(\textit{path},\textit{length})\leftarrow\textsc{ParseShortestPath}(\textit{answer})$
@@
else if $\textit{length}=\textit{ground\_truth}$ and path is valid in $G$ then